OpenStudy (anonymous):

How much money would you need to invest in an account with 7% interest compounded continuously if you want to have $5000 at the end of 5 years? (Round to the nearest cent.)

OpenStudy (anonymous):

do you know the formula for continuous compounding?

OpenStudy (anonymous):

Pe^rt

OpenStudy (anonymous):

right so, plug in the constants you know, what do you get?

OpenStudy (anonymous):

.35

OpenStudy (anonymous):

i don't think that's right, how did you get that number

OpenStudy (anonymous):

well i did.07*5 and got .35 then i put that in with e and got 1.41906

OpenStudy (anonymous):

okay you're almost there, what about the other constants

OpenStudy (anonymous):

what other constant? P ?

OpenStudy (anonymous):

yes, and what's it all equal to?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

IniitalPrincipal * e^(rt) = FinalValue You already computed the e^(rt), which was the hard part, now you have two variables left, one known and one unknown.

OpenStudy (anonymous):

3524e^.53

OpenStudy (anonymous):

3524

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yep! InitialPrincipal is P, and your desired final value is 5000 So P * e^(rt) = 5000 P * 1.42 = 5000 P = 3523.44